OpenStudy (anonymous):

If g(x) = x2 - 3, find g(4).

OpenStudy (ajprincess):

apply 4 for x.

OpenStudy (anonymous):

g(x)=x2-3 if x=4........ g(4)=4^2-3=16-3=13

OpenStudy (anonymous):

thnks